### v1

# Exploration of Quantum Computing in Materials Discovery for Direct Air Capture Applications

Abstract
Direct air capture is a promising method for mitigating climate change. Solid sorbents, such as metal-organic frameworks, have been considered for this task, but their potential for deployment at scale has not been fully realized. The computational discovery of sorbent materials is a daunting task, given the vast search space, and the fact that their real-world performance depends on their ability to selectively bind CO2 molecules while avoiding other more abundant flue gas components. Quantum computing can potentially accelerate the discovery of novel materials for direct air capture as an alternative way to compute binding energies. In this work, we demonstrate methods and algorithms that enable these calculations in current quantum computing devices. We simulate the potential energy surfaces of CO2, N2, and H2O molecules at the Mg+2metal center that represents the binding sites of typical metal-organic frameworks. We apply the qubit-ADAPT-VQE technique to run simulations on both classical and quantum hardware, and achieve reasonable accuracy while maintaining hardware efficiency, even when compared to more established methods such as UCCSD-VQE.
[…]
Centro Brasileiro de Pesquisas Físicas, Rio de Janeiro, 22290-180, RJ, Brazil
1 Introduction
Technologies for capturing greenhouse gases (GHG) are essential climate mitigation tools that can help us reduce GHG atmospheric concentration and limit global warming.1 Existing energy generation infrastructure, such as natural gas and coal-fired power plants, can be retrofitted to perform post-combustion capture of carbon dioxide. This strategy can extend the lifetime of existing infrastructure and minimize the impact of energy generation to our global carbon budget. Nevertheless, merely bringing emissions to zero is not sufficient for achieving our climate goals. We need to actively remove carbon dioxide from the atmosphere if we want to limit the temperature increase to manageable levels. Direct air capture (DAC) is an emerging technology that employs sorbents to capture CO2 from the atmosphere.
Liquid sorbent technologies, such as amine scrubbing, are the most well-established for capturing carbon dioxide from the air. Nevertheless, despite its fast adsorption rate and high CO2 saturation, amines present disadvantages that hinder their widespread implementation: they are corrosive, and due to their high desorption energy requirement, are responsible for 60-80% of total operating costs of CCS.2 Solid sorbent materials for carbon capture offer an alternative approach to overcome the limitations of amine-based (liquid) adsorption. This material class includes a diverse range of nanoporous solids, such as zeolites, metal-organic frameworks, covalent-organic frameworks, zeolitic imidazolate frameworks and porous polymer networks, among others. Porous solid sorbents can selectively adsorb CO2 without bond formation (i.e., via physisorption), which drastically lowers the regeneration energy and, therefore, the operating costs. Another advantage over amines is that the amount of CO2 adsorbed is not proportional to the amount of solid material, given that the adsorption takes place at the void space (e.g., cages, channels, or slits) that exists inside the solid. On the other hand, solid sorbents can be very sensitive to the presence of water vapor in the flue gas, leading to decreased selectivity due to competitive adsorption with CO2 and incomplete regeneration due to water passivation of adsorption sites.
Among the many solid materials classes considered, Metal-Organic Frameworks (MOFs) have attracted considerable attention due to their customizable porous structure and chemical diversity.3 They are composed of two types of building blocks – i.e., organic linkers and metal ions – which can be combined to form nanoporous structures with channels whose diameters range from a few to tens of Angstroms. The chemical and geometrical aspects can be leveraged independently to optimize both the absolute (adsorption capacity) and relative (adsorption selectivity) metrics that determine what is a good carbon capture material. The ample variety of MOFs, one of its main advantages, is also the origin of significant computational challenges for the selection of the best candidate material for a given application. There are databases of hypothetical MOFs,4, 5 as well as experimentally verified MOFs,6 amounting to $10^{5}-10^{6}$ materials in total.
(a) Conventional unit cell with 162 atoms
(b) Primitive unit cell with 54 atoms.
Figure 1: Atomistic visualisation of Mg-MOF-74 structure
In this work, we perform Quantum Chemistry calculations, on classical and on quantum hardware, to estimate the binding potential of a few molecules of interest present in the atmosphere (CO2, H2O and N2) to simplified model systems for carbon capture. We simplified the search space for a carbon capture material by focusing on the X-MOF-74 family (where X=Mg+2, Ni+2, Fe+2, Co+2, Zn+2, Mn+2, Cu+2) as a template for exploring the chemical diversity in MOFs. The MOF-74 family is known for the presence of coordinatively unsaturated metal ions (open metal sites) that can form a bond with adsorbate molecules, greatly influencing the adsorption strength and selectivity, especially at low pressures. The formation of chemical bonds with the adsorbate molecule is a phenomenon known as chemisorption and determines, to a great extent, the desorption energy cost.
1.1 Methods
Since the Mg-MOF-74 primitive cell, as shown in Figure 1, contains more than 50 atoms, we simplify the system even further by considering a single metal ion representing an open metal site and interacting with the adsorbate molecules. We do so because of the current limitations of quantum hardware which limits the width and depth of quantum circuits that can be reliably executed. We calculate Potential Energy Surfaces (PES) for several metal/molecule combinations and extract the equilibrium distance and binding energy for the systems represented in Figure 2. This proof-of-concept illustrates how one can use quantum computing to address the challenge of discovering materials for direct air capture.
(a) $M$ + H2O
(b) $M$ + N2
(c) $M$ + CO2
Figure 2: Molecular geometries and bond axes used in the PES calculations for (a) H2O, (b) N2 and (c) CO2 molecules in the vicinity of a metal ion $M$. For H2O, the ion is moved along the C2 rotation axis. For N2, the ion is moved along the N-N axis. For CO2, the ion is moved along the O-C-O axis
1.1.1 Species selection
Given the relevance of these metal ions to the application in MOFs, we chose to focus on Mg+2metal ion and its pairings with all the three provided gas molecules (H2O, CO2, and N2), amounting to six distinct systems. In short, Mg+2 + H2O should be the simplest system as it contains the least number of electrons.
1.1.2 Simulation methods
We chose PySCF,7, 8 as the driver for performing classical benchmarks. We ran three types of classical benchmarks with PySCF: Hartree-Fock (HF) and Density Functional Theory (DFT). The chosen basis set was 6-31g* and was kept consistent across all calculations, including for the quantum calculations. For DFT we chose the B3LYP functional. This particular choice of parameters was motivated by the methodology employed to compute 3D structures (SD-file) of the H2O, CO2, and N2 molecules found in the NIST Chemistry WebBook, from which we also took the molecular structures used in the calculations. 9
1.1.3 Considerations for Quantum devices
We chose Qiskit10 as the driver for performing quantum simulations using the Variational Quantum Eigensolver (VQE).11, 12, 13 We employed three methodological variants:
  • A usual VQE calculation with an application-specific, compute-heavy UCCSD ansatz,14 to be treated as the quantum benchmark, as it provides the best possible result for this kind of application. Unfortunately, it is challenging to use this approach on current quantum devices due to its complexity.
  • A usual VQE calculation with a general-purpose, hardware-efficient heuristic ansatz that can run on current quantum devices. In our case, we determined that an ansatz with RY and RZ rotations, which we call RYRZ, converges to the intended result despite requiring more iterations.
  • An implementation of the ADAPT-VQE algorithm15 that utilizes qubit operators in the operator pool. This method is called qubit-ADAPT-VQE and has the advantage of giving similar results to UCCSD while having much lower requirements on the quantum hardware.16
1.1.4 Active Space Selection
To limit the computational cost for the VQE calculations, we performed a suitable active space selection. We used a known method implemented within PySCF called atomic valence active space (AVAS).17 AVAS is an automated method to construct an active space for electronic structure calculations in strongly correlated chemical systems, where single-determinant wave functions are inadequate. It focuses on identifying specific valence atomic orbitals, such as transition metal d orbitals or those involved in bond dissociation, as a basis for defining the active space. The approach leverages density matrix embedding theory (DMET) to determine the minimal set of orbitals required to capture the essential electronic features, aiming for an active space no more than twice the size of the initially chosen orbitals. This method does not differentiate between “fragment” and “bath” orbitals, instead maintaining the occupied and virtual character of active orbitals, facilitating a natural truncation process to minimize active space size while capturing critical determinants.
2 Results and discussion
2.1 Convergence studies
We begin by evaluating the performance of qubit-ADAPT-VQE against the more standard VQE approaches. In Figure 3, we compare all these VQE simulations in terms of their convergence and accuracy with respect to the ground state energy value obtained from the direct diagonalization of the Hamiltonian using NumPy. We used the Mg+2 + H2O model system at 1.9 Å separation and an active space of six electrons and seven molecular orbitals (6,7). In Table 1, we display the runtime and total number of iterations needed for the convergence of each VQE ansatz. The qubit-ADAPT-VQE took 3x and 64x fewer iterations, respectively, than the UCCSD-VQE and RYRZ-VQE ansatze. In terms of run time, the qubit-ADAPT-VQE was 5x and 12x faster than the UCCSD-VQE and RYRZ-VQE ansatze, respectively.
Figure 3: Convergence to the ground state energy calculated by UCCSD-VQE (green), RYRZ-VQE (red) and qubit-ADAPT-VQE (blue) as a function of number of iterations. The result of the direct diagonalization of the fermionic Hamiltonian using NumPy (black) is shown as the target. All ansatze converge, to within chemical accuracy ($\sim 1\ \text{mHa}$), of the target result. However, the qubit-ADAPT-VQE ansatz reaches convergence in considerably fewer iterations than the UCCSD-VQE and RYRZ-VQE ansatze. The energy values correspond to the computed part in the (6,7) active space.
Table 1: The qubit-ADAPT-VQE ansatz converges in 3x and 116x fewer iterations, and is 5x and 25x faster than the UCCSD-VQE and RYRZ-VQE ansatze, respectively.
UCCSD-VQE RYRZ-VQE qubit-ADAPT-VQE
Run time (s) 25 151 6
Total # of iterations 133 4892 42
Energy Value (Ha) -56.93440792 -56.93383614 -56.93411960
The qubit-ADAPT-VQE ansatz accomplishes fast convergence while remaining hardware efficient, as we can see from Table 2 comparing the sizes of the circuits between methods with the largest circuit that qubit-ADAPT-VQE uses. These circuit sizes are expressed in terms of the gates that IBM quantum devices execute and were built using the transpilation procedure from Qiskit. The most important metric here is the number of CNOT gates as those have error rates that are orders of magnitude higher than the single-qubit gates.
Table 2: Comparing the circuit sizes generated by different ansatze. qubit-ADAPT-VQE manages to get satisfactory results while using the shallowest circuits. Circuits are described in terms of the number of gate operations and depth required to run on IBM quantum devices.
UCCSD-VQE RYRZ-VQE qubit-ADAPT-VQE
# of CNOT gates 1610 27 20
# of Rz gates 1191 125 37
# of Sx gates 878 80 19
# of X gates 44 5 1
Depth 3255 35 30
The performance advantage of qubit-ADAPT-VQE comes from the fact that the ansatz grows on demand from an operator pool as the convergence thresholds require. We created the operator pool by first building an UCSSD ansatz and then removing Pauli strings with an even number of Y operations.16 This way we are left only with the “odd” Pauli Strings which are the ones relevant for the energy calculation. We chose the SLSQP optimizer to run the simulations. Finally, we defined the convergence thresholds: maximum number of iterations as $4$, maximum operator gradient as $10^{-4}$, and maximum change of the energy expectation value between consecutive iterations as $10^{-3}$. These convergence thresholds ensure chemical accuracy ($\sim 1\ \text{mHa}$) with respect to the UCCSD-VQE and NumPy results. These settings were used for all qubit-ADAPT-VQE simulations.
We did not include a noise model in the simulations, as we are mostly interested in comparing how our ansatze and methods perform against a benchmark provided by the direct diagonalization of the Hamiltonian with NumPy on a classical computer. Only after their convergence and accuracy are well established, we can repeat the best procedure on the quantum hardware.
We had to choose a suitable active space to run these quantum simulations in a reasonable time. Whenever possible, we utilized the AVAS automated selection method mentioned before, but if the given active space required more than 15 qubits, we would default that calculation to an active space of six electrons and seven molecular orbitals (6,7) chosen around the HOMO-LUMO level with orbitals ordered by energy contribution. This active space requires 10-11 qubits, which we adopted as a reasonable compromise between quality and runtime. All VQE and PySCF simulations were executed on a laptop (Apple MacBook Air M1 2020). We ran the simulations on a slim environment so that anyone with access to our code repository could easily reproduce our results.18
2.2 Mg+2 systems
Due to the excellent cost-benefit obtained from the qubit-ADAPT-VQE ansatz in the simulations, we selected this as our best candidate to run experiments on the quantum hardware. However, to run these calculations on the available quantum hardware, we had to further reduce our active space to three electrons and four molecular orbitals (3,4), which only required 4 qubits. Despite nominally supporting higher qubit counts, current devices are not suitable for reliably running circuits of the required depth for this application due to a lack of fault tolerance. Again, the molecular orbitals were chosen around the HOMO-LUMO level and ordered by energy. We set the limit on the number of shots to 1024, which takes from five to eight hours depending on the separation distance.
The 4-qubit active space limitation led us to focus only on the Mg+2 systems for hardware experiments. This active space corresponds to $\approx 5\%$ of the energy contributions to the ground state, therefore we saw no point in simulating systems with a higher electron count as it would mean an even lower representation of the results obtained. We used Qiskit Runtime and ibmq_montreal to run all our quantum hardware experiments, and we employed error mitigation techniques such as Twirled Readout Error Extinction (T-REx).19.
The PES was built by computing the energy for different separations between metal ions and molecules along the bond axis. For all simulations and experiments, we evaluate the energy at distances between 0.3 and 3.6 Å, in increments of 0.3 Å.
After we calculate the energies, we fit a Morse potential20 to the data points to help us calculate precise equilibrium distances and molecular binding energies. This potential is a suitable inter-atomic model for the energy of diatomic molecules. In the PES plots, we offset the energies so that the asymptotic limit at infinite separation coincides with the zero-energy point to facilitate the extraction of the binding energy.
We simulated Mg+2 binding with all of the suggested gas molecules using the three VQE methods mentioned before: RYRZ-VQE, UCCSD-VQE, and qubit-ADAPT-VQE, on a (6,7) active space. We evaluate their accuracy by comparing their PES to that obtained via direct diagonalization of the Hamiltonian with NumPy.
In Figure 4, we observe that qubit-ADAPT-VQE and UCCSD-VQE successfully achieve the NumPy result, while RYRZ-VQE only matches effectively the simplest system of the three, Mg+2 + H2O. The RYRZ-VQE results for the N2 and CO2 systems overestimate the energy to the left of the minimum and underestimate to the right, affecting the Morse potential fit and leading to different observables.
(a) Mg+2 + H2O
(b) Mg+2 + N2
(c) Mg+2 + CO2
Figure 4: UCCSD-VQE (green circles) and qubit-ADAPT-VQE (blue circles) match the NumPy (magenta triangles) benchmark in all three cases, however, RYRZ-VQE (red circles) only matches the benchmark for the simplest system, namely, Mg+2 + H2O.
Based on the accuracy of qubit-ADAPT-VQE results in Figure 4, and the superior performance over UCCSD-VQE demonstrated in Figure 3, Table 1 and Table 2, we confirm the choice of qubit-ADAPT-VQE for running VQE on quantum hardware. We limit ourselves to the smallest possible system, namely, Mg+2 + H2O for the reasons explained previously. We ran qubit-ADAPT-VQE on quantum hardware with an active space of (3,4) to comply with the limitations of current quantum devices.
The results, shown in Figure 5, also provide a comparison to a quantum hardware calculation. The Morse potential fit to the hardware result matches quite closely those of the direct diagonalization via NumPy and simulated VQE, despite the level of noise that is inherent to calculations on current quantum devices. In this hardware experiment, we used 1024 shots and employed the T-REx noise mitigation technique which proved sufficient to handle the noise profile of the quantum device. Increasing the number of shots did not improve the accuracy of the experimental values, and using more robust error mitigation strategies, such as Zero Noise Extrapolation, would increase the already large variance in the results.
The plots show the calculated energies at selected distances represented by solid symbols and the Morse potential fit represented by solid lines. In the following plots, circles represent VQE simulations, triangles represent the direct diagonalization with NumPy, squares represent the PySCF calculations and diamonds represent the hardware experiments.
Figure 5: Potential Energy Surfaces obtained from running VQE on quantum hardware (black diamonds) compared to a qubit-ADAPT-VQE simulation (blue circles) and NumPy direct diagonalization (purple triangles). The adjusted Morse potentials exhibit a considerable overlap, despite the level of noise present in the hardware data points, which leads to a reasonable agreement between the observables shown in Table 3.
We extracted observables such as the equilibrium distance and the binding energy from the adjusted Morse potentials to the Mg+2 + H2O PES curves. Although we cannot claim to have achieved agreement to within chemical accuracy ($\sim 1\ \text{mHa}$), on Table 3 we observe an excellent match between qubit-ADAPT-VQE and NumPy, and a reasonably good match with ibmq_montreal, considering the higher uncertainties caused by the quantum device noise.
Table 3: Comparison of Equilibrium Distance and Binding Energy for Mg+2 + H2O obtained from the fitted Morse potential to results obtained with different methods. All methods are in agreement, to some extent, when we consider the uncertainties involved in the fitting procedure and the inherent noise of the quantum hardware data points.
ibmq_montreal qubit-ADAPT-VQE NumPy
Equilibrium Distance (Å) $1.8\pm 0.3$ $1.87\pm 0.01$ $1.87\pm 0.01$
Binding energy (Ha) $-0.1\pm 0.4$ $-0.1\pm 0.02$ $-0.1\pm 0.02$
We compare the potential energy surfaces computed using classical methods from PySCF, such as HF and DFT, to the best-in-class simulated VQE method, namely, qubit-ADAPT-VQE in a (6,7) active space. In Figure 6, we show the agreement between quantum methods and their classical benchmarks. In particular, for Mg+2 + H2O, we also include in Figure 6(a) the comparison to the quantum hardware experiments running qubit-ADAPT-VQE with ibmq_montreal in a (3,4) active space.
Overall, the simulated qubit-ADAPT-VQE results match very well the HF benchmark for all Mg+2-based systems studied, with the DFT benchmark coming quite close. The PES obtained by running in the quantum hardware, despite its inherent noise, is also in good agreement with the classical benchmarks. This conclusion is supported not only by the visual inspection of Figure 6 but also by a comparison of the numeric results of the observables extracted from the PES fit in Table 4. Here, the Binding Energy values all agree within the numeric uncertainty, and while the equilibrium distances between qubit-ADAPT-VQE and HF concur, they are found to be 2-3% higher than what is predicted by DFT.
It’s worth noting the high uncertainties in the hardware results. This comes as a cost of using error mitigation techniques. while accuracy is increased there is a loss in precision. Considering its already relatively high value we opted not to use more robust error mitigation techniques that would increase it further.
(a) Mg+2 + H2O
(b) Mg+2 + N2
(c) Mg+2 + CO2
Figure 6: Potential energy surfaces for the Mg+2 systems obtained from VQE experiments on quantum hardware using (3,4) active space (black diamonds) and VQE simulations in a (6,7) active space (blue circles) compared to PySCF-based classical benchmarks using the HF (red squares) and DFT (green squares) methods. The simulated qubit-ADAPT-VQE result matches the HF benchmark for the (a) H2O, (b) N2 and (c) CO2 molecules. The result from the hardware experiments, despite their inherent noise, also show accurate values within 0.1 Angstrom of both the simulated VQE and the classical benchmarks.
Table 4: Comparison of Equilibrium Distances and Binding Energies for Mg+2-based systems obtained from the fitted Morse potential across different methods, both quantum (simulations and experiments) and classical. The Binding Energy values all agree to within the numeric uncertainty. The equilibrium distances agree between qubit-ADAPT-VQE and HF, but are 2-3% higher than what is predicted by DFT.
$Mg^{+2}+H_{2}O$
ibmq_montreal qubit-ADAPT-VQE PySCF HF PySCF DFT
Equilibrium Distance (Å) $1.8\pm 0.3$ $1.87\pm 0.01$ $1.868\pm 0.004$ $1.83\pm 0.02$
Binding energy (Ha) $-0.1\pm 0.4$ $-0.10\pm 0.02$ $-0.10\pm 0.01$ $-0.1\pm 0.02$
$Mg^{+2}+N_{2}$
ibmq_montreal qubit-ADAPT-VQE PySCF HF PySCF DFT
Equilibrium Distance (Å) - $1.95\pm 0.04$ $1.96\pm 0.05$ $1.90\pm 0.07$
Binding energy (Ha) - $-0.10\pm 0.04$ $-0.10\pm 0.06$ $-0.1\pm 0.08$
$Mg^{+2}+CO_{2}$
ibmq_montreal qubit-ADAPT-VQE PySCF HF PySCF DFT
Equilibrium Distance (Å) - $1.86\pm 0.02$ $1.86\pm 0.02$ $1.81\pm 0.02$
Binding energy (Ha) - $-0.10\pm 0.02$ $-0.10\pm 0.02$ $-0.1\pm 0.02$
The experimental binding energies for CO2 on Mg-MOF-74 are in the range of 39-47 kJ/mol.21 These values translate to 15-18 mHa per molecule, which is about 6x smaller than the results we obtained. This illustrates the limits of the oversimplification made by reducing the Mg-MOF-74 structure to a single metal ion.
2.3 Conclusions
As a key result of our investigation, the qubit-ADPAT-VQE achieves chemical accuracy when compared to UCCSD-VQE while maintaining hardware efficiency. This means it can give us an advantage when using quantum hardware by having a problem-specific ansatz that is also hardware efficient. Note that, despite the noisy results, we can fit the hardware data with a Morse potential so as to agree with Hartree-Fock predictions and the noiseless simulations.
Our results could be improved with more robust noise mitigation strategies, such as Zero Noise Extrapolation. It would also be of interest to explore larger systems and active spaces in order to improve our preliminary results. Although for that first we need to explore better techniques for active space selection.
Throughout our simulations, for Mg+2-based systems, we observed less satisfactory results when simulating the adsorption of N2 molecules. Both the accuracy and the precision deteriorated across all methods used when compared to the other molecules. As a potential explanation, we speculate that the triple bond, which involves six electrons at once, is particularly hard to model. When the positively charged metal ion approaches, it perturbs at once all six active electrons involved in the triple bond. This seems to be especially problematic for the calculations limited to an active space, since the (6,7) active space is entirely consumed by these six electrons, leaving many other relevant electrons out of the calculation.
We also observed a systematic overestimation of the HF equilibrium distance with respect to the one obtained by DFT. This discrepancy was less than 5% and may be attributed to the presence of exchange-correlation energy contributions that are absent in HF. This requires further investigation.
We note that current hardware limitations are mainly due to reliability, such as resistance to noise, and computation speed, meaning the number of circuits the hardware can run at a given time. The evolution of the available quantum volume and circuit-layer operations that can be performed at a given time interval will soon open pathways for larger-scale computation of CO2 capture in MOFs for DAC applications.
{acknowledgement}
The authors thank Max Rossmannek (IBM Research – Zurich) for providing help in implementing qubit-ADAPT-VQE and the broader Qiskit community for giving support via the Slack channel and GitHub pages. We thank Vidushi Sharma (IBM Research – Almaden) for her support in the interpretation of the classical benchmark methods. We thank Felipe Lopes de Oliveira (IBM Research – Brazil) for suggesting the Morse potential fit as a means to extract the observables. We thank Ivan S. Oliveira (CBPF) for insightful discussions on quantum computing. The authors would like to acknowledge the support of Alexandre Pfeifer and Bruno Flach (IBM Research - Brazil) for the realisation of this work.
3 Code availability
The code used to generate the results presented in this paper is available in a public GitHub repository.18 This repository also includes additional experimental results that were not detailed in the manuscript. For ease of use, the repository contains a README file that explains how the additional results complement those discussed in the paper and guides the navigation of the supplementary material. Should there be any issues or questions regarding the code and its implementation, we encourage readers to reach out.